OpenStudy (anonymous):

Solve: 4 3x - 1 = 8 x + 1

OpenStudy (anonymous):

Is that 43x? If it is then: 43x-8x=1-1 35x=0 x=0/35 = 0

OpenStudy (anonymous):

yes it is thank you :)